### v1

# thanks: Corresponding author

Abstract
In recent years, jumps in the particle circular-orbit radius $r_{\rm c}$ have been proposed as observational signatures of black hole phase transitions, supported by the general consensus that $r_{\rm c}$ increases monotonically with the horizon radius $r_{h}$. However, in this paper, we will show that this monotonicity is not universally valid. By deriving a universal, ensemble-independent criterion for $\mathrm{d}r_{\rm c}/\mathrm{d}r_{h}$ and applying it to the $d$-dimensional Reissner-Nordström-Anti-de Sitter black hole, we reveal a sharp dichotomy: while $r_{\rm c}(r_{h})$ exhibits strictly monotonicity along isobars, it typically becomes non-monotonic along isotherms (except for the photon sphere of Schwarzschild-Anti-de Sitter black hole)–a behavior that has been overlooked in previous studies. Nevertheless, we rigorously prove that within the phase-transition region itself, the monotonicity of $r_{\rm c}(r_{h})$ is preserved. However, the generalizability of this result to other spacetimes requires case-by-case verification, and its validity could have notable implications for the observability of the jump in $r_{\rm c}$. Interestingly, non-monotonicity itself acts as a new observational probe: the monotonicity contrast in $r_{\rm c}(r_{h})$ distinguishes isobaric from isothermal evolution, providing a complementary thermodynamic probe. Furthermore, we show that the critical exponents extracted from $r_{\rm c}$ match those obtained from $r_{h}$, independent of ensemble. Remarkably, all these phenomena are closely tied to the thermodynamic first law.
I Introduction
Black hole thermodynamics has evolved into a central theme of gravitational physics. The Hawking-Page transition first revealed a first-order transition between thermal Anti-de Sitter (AdS) radiation and large AdS black holes [1]. Subsequent work has uncovered thermodynamic phase transitions in various black-hole spacetimes, including Van der Waals-type small/large transitions in AdS black holes [2, 3, 4, 5, 6, 7], thereby establishing a rich thermodynamic phase structure.
Yet direct observation of these thermodynamic phenomena remains limited. In contrast, optical and dynamical features–such as the photon sphere and the innermost stable circular orbit (ISCO)–are directly observable and encode essential geometric information relevant to shadows and accretion physics [8, 9, 10]. The EHT image of M87* made this explicit, revealing an emission ring whose boundary traces the photon orbit [11].
This naturally raises the possibility that black hole thermodynamics could be probed through optical features. Compelling evidence comes from the work of Wei et al., who showed that in charged AdS black holes, both the photon-sphere radius and the minimum impact parameter exhibit Van derWaals-type non-monotonicity and jump discontinuities, behaving as mean-field order parameters [12]. Subsequently, related signatures involving photon orbits and ISCOs have been identified across a variety of spacetimes and modified gravitational theories [13, 14, 15].
A central consideration in these studies is the relationship between the circular-orbit radius $r_{\rm c}$ and the horizon radius $r_{h}$. Across a first-order small/large transition, $r_{h}$ undergoes a discontinuous jumps, and the monotonicity $r_{\rm c}(r_{h})$ ensures that $r_{\rm c}$ also jumps accordingly–making it a natural geometric order parameter (see Appendix A for a brief mathematical derivation). This expectation was later formalized by Zhang et al., who proved monotonic growth of both radii (the photon-sphere radius and ISCO radius) with $r_{h}$ in general static spherical geometries–a result that has been widely applied in subsequent work [16].
However, we find that once more than one thermodynamic parameter is allowed to vary, this monotonicity claim fails. To illustrate this, we present a simple counterexample based on the Schwarzschild-AdS (Sch-AdS) black hole: the photon-sphere radius $r_{\rm ps}=3M$ depends only on the mass $M$, whereas the horizon radius $r_{h}$ depends on both $M$ and $P$, where $P=-\Lambda/(8\pi)$ with $\Lambda$ being the negative cosmological constant. Increasing $P$ decreases $r_{h}$, with $r_{\rm ps}$ unchanged. Consequently, there exist smooth paths in $(M,P)$ parameter space along which $r_{h}$ increases and $r_{\rm ps}$ decreases. Figure 1 provides an explicit example, and Appendix B presents a fully analytic realization. This breakdown of monotonicity in $r_{\rm c}(r_{h})$ has been neglected in previous studies. The reason is straightforward: most analyses have been restricted to the isobaric ensemble, where $P$ is held fixed and only $M$ varies, resulting in both $r_{\rm c}$ and $r_{h}$ typically grow monotonically with $M$. Even in studies that employed the isothermal ensembles (with both $M$ and $P$ varying), the monotonicity of $r_{\rm c}(r_{h})$ was not systematically examined, leading to the omission of non-monotonic behavior. Thus, a genuine understanding of how circular orbits encode thermodynamic information requires analyzing the monotonicity of $r_{\rm c}(r_{h})$ along arbitrary parameter paths–precisely the general framework developed in this work.
Figure 1: Behavior of $M(r_{h},P)$ in the Sch-AdS spacetime. Thin black solid curves denote contours of constant mass $M$. Along the dashed trajectory from A to B, the horizon radius $r_{h}$ increases while both $M$ and the photon-sphere radius $r_{\rm ps}=3M$ decrease, showing that $r_{\rm ps}(r_{h})$ is strictly monotonically decreasing along this path.
Our strategy is as follows. In Sec. II, we review the conditions for photon spheres and ISCOs in a general static, spherically symmetric spacetime, and derive a unified expression for $\mathrm{d}r_{\rm c}/\mathrm{d}r_{h}$ along an arbitrary parameter path. In Sec. III, we analyze the monotonicity of $r_{\rm c}(r_{h})$ for $d$-dimensional Reissner-Nordström-AdS (RN-Ads) black holes, focusing on the isobaric ($T$-$S$) and isothermal ($P$-$V$) thermodynamic paths. In Sec. IV, we show that the critical exponents remain invariant under the reparametrization $r_{h}\!\mapsto r_{\rm c}$, demonstrating that $r_{\rm c}$ defines the same universality class as $r_{h}$. Finally, Sec. V summarizes and discusses our main results.
II Monotonicity criterion
II.1 Review of circular orbits
Consider a static, spherically symmetric spacetime with the line element
$\mathrm{d}s^{2}=-f(r;\lambda)\,\mathrm{d}t^{2}+\frac{\mathrm{d}r^{2}}{f(r;\lambda)}+r^{2}\mathrm{d}\Omega^{2},$ (1)
where $f(r;\lambda)$ is a smooth function of the radial coordinate $r$, while $\lambda=(\lambda_{1},\ldots,\lambda_{n})$ denotes the set of black-hole parameters, such as mass $M$, charge $q$, and so on. The Lagrangian of a particle moving in the spacetime determined by Eq. (1) reads
$\mathcal{L}=\tfrac{1}{2}\,g_{\mu\nu}\dot{x}^{\mu}\dot{x}^{\nu}=-\tfrac{1}{2}\,\epsilon,\quad\epsilon=\begin{cases}0,&\text{massless particle (photon)},\\ 1,&\text{massive particle}.\end{cases}$ (2)
Where the overdot denotes differentiation with respect to the affine parameter. Without loss of generality, we set $\theta=\pi/2$, restricting the motion to the equatorial plane. Then, the Lagrangian reduces to
$\mathcal{L}=\frac{1}{2}\!\left[-f(r;\lambda)\,\dot{t}^{\,2}+\frac{\dot{r}^{\,2}}{f(r;\lambda)}+r^{2}\dot{\phi}^{\,2}\right]=-\tfrac{1}{2}\,\epsilon.$ (3)
Since the metric is stationary and axisymmetric, the corresponding cyclic coordinates $t$ and $\phi$ give rise to two conserved quantities,
$E:=-\frac{\partial\mathcal{L}}{\partial\dot{t}}=f(r;\lambda)\,\dot{t},\qquad L:=\frac{\partial\mathcal{L}}{\partial\dot{\phi}}=r^{2}\dot{\phi},$ (4)
which represent the particle’s energy and angular momentum, respectively. Substituting Eq. (4) into Eq. (3) yields the following radial equation of motion
$\dot{r}^{\,2}=E^{2}-f(r)\!\left(\epsilon+\frac{L^{2}}{r^{2}}\right)\equiv\mathcal{R}(r).$ (5)
The particle’s circular orbits at $r=r_{c}$ are determined by
$\mathcal{R}(r_{c})=0,\qquad\mathcal{R}^{\prime}(r_{c})=0.$ (6)
Solving Eq. (6) for a massless particle with $\epsilon=0$ gives
$\Phi_{0}(r;\lambda):=r\,\partial_{r}f(r;\lambda)-2f(r;\lambda)=0,\quad\text{at}\,r=r_{\rm ps},$ (7)
where $r_{\rm ps}$ denotes the photon-sphere radius. The condition for an unstable photon sphere further requires
$\mathcal{R}^{\prime\prime}(r_{\rm ps})<0,$ (8)
which is equivalent to
$\Gamma_{0}(\lambda):=\left.\partial_{r}\Phi_{0}(r;\lambda)\right|_{r_{\rm ps}}<0.$ (9)
The ISCO of a massive particle with $\epsilon=1$ marks the transition between stable and unstable circular motion and is defined by the marginal stability condition
$\mathcal{R}(r_{\rm isco})=\mathcal{R}^{\prime}(r_{\rm isco})=\mathcal{R}^{\prime\prime}(r_{\rm isco})=0,$ (10)
which can be recast into a universal form,
$\Phi_{1}(r;\lambda)$ $=r\,f(r;\lambda)\,\partial_{r}^{2}f(r;\lambda)-2r\big(\partial_{r}f(r;\lambda)\big)^{2}$ (11)
$+3f(r;\lambda)\,\partial_{r}f(r;\lambda)=0,\quad\text{at}\,r=r_{\rm isco}.$
To ensure that circular orbits are stable outside ISCO but unstable inside, one further requires its radius $r_{\rm isco}$ to satisfy
$\Gamma_{1}(\lambda):=\left.\partial_{r}\Phi_{1}(r;\lambda)\right|_{r_{\rm isco}}>0.$ (12)
Summarizing the above conditions, we have
$\Phi_{\epsilon}(r;\lambda)=0,\quad\text{at}\,r=r_{c}(\epsilon;\lambda):=\begin{cases}r_{\rm ps}(\lambda),&\epsilon=0,\\ r_{\rm isco}(\lambda),&\epsilon=1,\end{cases}$ (13)
and the corresponding stability criteria
$\Gamma_{\epsilon}(\lambda)\begin{cases}<0,&\epsilon=0,\\[4.0pt] >0,&\epsilon=1.\end{cases}$ (14)
II.2 Derivation of the Monotonicity Criterion
To investigate the monotonic relation between the circular-orbit radius $r_{c}$ and the event-horizon radius $r_{h}$, we first express them in terms of the previously defined functions $\Phi_{\epsilon}(r;\lambda)$ and $f(r;\lambda)$
$\Phi_{\epsilon}(r_{c};\lambda)=0,\qquad f(r_{h};\lambda)=0.$ (15)
According to the implicit function theorem, if
$\left.\partial_{r}\Phi_{\epsilon}\right|_{(r_{\mathrm{c}};\lambda)}\neq 0,\qquad\left.\partial_{r}f\right|_{(r_{h};\lambda)}\neq 0,$ (16)
then there exist locally smooth functions $r_{\mathrm{c}}(\lambda)$ and $r_{h}(\lambda)$ in the neighborhoods of $(r_{\mathrm{c}};\lambda)$ and $(r_{h};\lambda)$, respectively. Fortunately, according to Eqs. (9) and (12), one finds
$\left.\partial_{r}\Phi_{\epsilon}\right|_{(r_{\mathrm{c}};\lambda)}=\Gamma_{\epsilon}\neq 0,$ (17)
and
$\left.\partial_{r}f\right|_{(r_{h};\lambda)}=2\kappa>0,$ (18)
where $\kappa=\tfrac{1}{2}\,\partial_{r}f(r;\lambda)\big|_{(r_{h};\lambda)}$ is the surface gravity of the black hole, which is positive for any nonextremal case. Hence, the conditions (16) are automatically satisfied.
Based on this, Eq. (15) can in principle be differentiated with respect to $\lambda$. Since the parameter set $\lambda$ generally contains several components, it is convenient to introduce a smooth path $\lambda(t)$ in the parameter space, along which all quantities become functions of a single real variable $t$. This allows us to directly apply the chain rule, yielding
$\frac{\mathrm{d}}{\mathrm{d}t}\,\Phi_{\epsilon}\big(r_{\mathrm{c}}(t);\lambda(t)\big)=\partial_{r_{\mathrm{c}}}\Phi_{\epsilon}\,\dot{r}_{\mathrm{c}}+\langle\nabla_{\lambda}\Phi_{\epsilon},\,\dot{\lambda}\rangle=0,$ (19)
$\frac{\mathrm{d}}{\mathrm{d}t}\,f\big(r_{h}(t);\lambda(t)\big)=\partial_{r_{h}}f\,\dot{r}_{h}+\langle\nabla_{\lambda}f,\,\dot{\lambda}\rangle=0,$ (20)
where the overdot denotes differentiation with respect to $t$, and $\langle\nabla_{\lambda}F,\dot{\lambda}\rangle\equiv\sum_{i=1}^{n}(\partial_{\lambda_{i}}F)\,\dot{\lambda}_{i}$. Combining Eqs. (19) and (20), we finally obtain
$\frac{\mathrm{d}r_{\mathrm{c}}}{\mathrm{d}r_{h}}=\frac{\dot{r}_{\mathrm{c}}}{\dot{r}_{h}}=\frac{2\kappa}{\Gamma_{\epsilon}}\,\frac{\langle\nabla_{\lambda}\Phi_{\epsilon},\,\dot{\lambda}\rangle}{\langle\nabla_{\lambda}f,\,\dot{\lambda}\rangle},\qquad(\dot{r}_{h}\neq 0),$ (21)
where we have written $\Phi_{\epsilon}$ and $f$ as shorthand for $\Phi_{\epsilon}(r_{\mathrm{c}};\lambda)$ and $f(r_{h};\lambda)$, respectively, and this convention will be used throughout the paper.
The criterion (21) shows that the monotonic relation between $r_{\mathrm{c}}$ and $r_{h}$ is governed by two independent factors. The first factor, $2\kappa/\Gamma_{\epsilon}$, is purely geometric, determined by the intrinsic properties of the black hole and its circular orbits. As discussed earlier, it remains nonvanishing and therefore does not affect the monotonicity of $r_{\mathrm{c}}(r_{h})$. The second factor, $\langle\nabla_{\lambda}\Phi_{\epsilon},\dot{\lambda}\rangle/\langle\nabla_{\lambda}f,\dot{\lambda}\rangle$, encodes how the parameters evolve along a chosen path in the thermodynamic space. It therefore determines whether the overall monotonicity of $r_{\mathrm{c}}(r_{h})$ is preserved or violated, emphasizing the importance of considering multiparameter constraints in black hole thermodynamics, as discussed in the Introduction.
In the next section, we apply the criterion (21) to two physically relevant constraints in black hole thermodynamics, namely the isobaric and isothermal processes, and investigate the monotonicity of $r_{\mathrm{c}}(r_{h})$ for the $d$-dimensional RN-AdS black hole.
III Examples of $d$-Dimensional Black Holes under Isobaric and Isothermal Constraints
In $d$-dimensional spacetime, a static and spherically symmetric AdS black hole is described by the line element
$\mathrm{d}s^{2}=-f(r)\,\mathrm{d}t^{2}+\frac{\mathrm{d}r^{2}}{f(r)}+r^{2}\,\mathrm{d}\Omega_{d-2}^{2},$ (22)
where $\mathrm{d}\Omega_{d-2}^{2}$ denotes the line element of a unit (d2)(d-2)-sphere. The metric function takes the general form
$f(r)=1-\frac{c_{d}M}{r^{d-3}}+\frac{q^{2}}{r^{2(d-3)}}+\frac{16\pi P}{(d-1)(d-2)}\,r^{2},$ (23)
where $M$ is the ADM mass, $q$ is the charge parameter, and $P=-\Lambda/(8\pi)$ represents the thermodynamic pressure associated with the cosmological constant $\Lambda$. The coefficients
$c_{d}=\frac{16\pi}{(d-2)\Omega_{d-2}},\qquad\Omega_{d-2}=\frac{2\pi^{(d-1)/2}}{\Gamma((d-1)/2)},$ (24)
depend only on the spacetime dimension, with $\Omega_{d-2}$ being the area of the unit (d2)(d-2)-sphere. The metric (23) reproduces several well-known limits:
  • $q=0$: the $d$-dimensional Sch-AdS black hole;
  • $q\neq 0$: the RN-AdS black hole;
  • $P\to 0$: the asymptotically flat Schwarzschild or RN solution in $d$ dimensions.
III.1 Isobaric Process
The isobaric process corresponds to the $T$-$S$ ensemble commonly employed in black hole thermodynamics. Here, all thermodynamic variables except the mass remain fixed. Thus, we may take $\lambda=(M;\lambda_{\mathrm{fixed}})$ and choose $M$ itself as the path parameter. This leads to $\dot{r}_{h}={\mathrm{d}r_{h}}/{\mathrm{d}M}$. To examine whether $\dot{r}_{h}$ can vanish, we consider the thermodynamic first law
$\mathrm{d}M=T\,\mathrm{d}S\quad\Rightarrow\quad\frac{\mathrm{d}r_{h}}{\mathrm{d}M}=\frac{1}{T\,S^{\prime}(r_{h})},$ (25)
where $T={2\kappa}/{4\pi}=\left.{\partial_{r}f(r;\lambda)}/{4\pi}\right|_{(r_{h};\lambda)}>0$ is the Hawking temperature, and $S^{\prime}(r_{h})=\mathrm{d}S/\mathrm{d}r_{h}$. In $d$-dimensional general relativity, S=A/4S={A}/{4} with $A=\Omega_{d-2}r_{h}^{\,d-2}>0$, so that
$S^{\prime}(r_{h})=\frac{\Omega_{d-2}}{4}\,(d-2)\,r_{h}^{\,d-3}>0\quad\Rightarrow\quad\frac{\mathrm{d}r_{h}}{\mathrm{d}M}>0.$ (26)
Hence, taking $M$ as the path parameter is fully justified, and the general criterion (21) simplifies to
$\frac{\mathrm{d}r_{\mathrm{c}}}{\mathrm{d}r_{h}}=-\,\frac{1}{\Gamma_{\epsilon}}\,\frac{\mathrm{d}M}{\mathrm{d}r_{h}}\,\partial_{M}\Phi_{\epsilon}.$ (27)
The monotonicity of $r_{\rm c}(r_{h})$ is therefore governed by $\partial_{M}\Phi_{\epsilon}$. Although no fundamental law requires $\partial_{M}\Phi_{\epsilon}\neq 0$, the mass $M$ determines the background geometry and hence controls the response of circular orbits, which generically vary with $M$.
For the metric (23), we obtain
$\partial_{M}\Phi_{0}$ $=(d-1)\frac{c_{d}}{r_{\rm ps}^{\,d-3}}>0,$ (28)
$\partial_{M}\Phi_{1}$ $=-\frac{2\,r_{\rm isco}^{\,1-4d}}{(d-1)(d-2)}\,\sqrt{h(r_{\rm isco},q,P,d)}<0,$
where the explicit expression of $h$ is omitted for brevity. Substituting Eq. (28) into Eq. (27) and using Eqs. (14) and (26), we find
$\epsilon=0:$ $\quad-\frac{1}{\Gamma_{0}}\frac{\mathrm{d}M}{\mathrm{d}r_{h}}>0,\ \partial_{M}\Phi_{0}>0\ \Rightarrow\ \frac{\mathrm{d}r_{\rm ps}}{\mathrm{d}r_{h}}>0,$ (29)
$\epsilon=1:$ $\quad-\frac{1}{\Gamma_{1}}\frac{\mathrm{d}M}{\mathrm{d}r_{h}}<0,\ \partial_{M}\Phi_{1}<0\ \Rightarrow\ \frac{\mathrm{d}r_{\rm isco}}{\mathrm{d}r_{h}}>0.$ (30)
Hence, along an isobaric process, both $r_{\rm ps}(r_{h})$ and $r_{\rm isco}(r_{h})$ increase monotonically. The opposite signs of $\Gamma_{0}$ and $\Gamma_{1}$ are exactly compensated by those of $\partial_{M}\Phi_{0}$ and $\partial_{M}\Phi_{1}$, resulting in a consistent monotonic relation between $r_{\rm c}$ and $r_{h}$. This implies that, during an isobaric process, a small-to-large black hole phase transition necessarily induces a corresponding discontinuous jump in both $r_{\mathrm{ps}}$ and $r_{\mathrm{isco}}$. Such abrupt changes could, in principle, be detected through astronomical observations, providing a potential signature of black hole phase transitions. These findings are consistent with the earlier results reported by Wei, Zhang, et al.
Next, we take $d=4$ as an explicit example and compute the expression of ${\mathrm{d}r_{\rm c}}/{\mathrm{d}r_{h}}$. For the four-dimensional (4D) RN-AdS black hole, one obtains
$\frac{\mathrm{d}M}{\mathrm{d}r_{h}}$ $=\frac{2\kappa}{{c_{d}}/{r_{h}^{\,d-3}}}=\frac{2(-{q^{2}}/{r_{h}^{3}}+{1}/{r_{h}}+8\pi P\,r_{h})}{{c_{d}}/{r_{h}^{\,d-3}}}>0,$
$\Gamma_{0}$ $=\frac{4q^{2}-2r_{\rm ps}^{2}}{r_{\rm ps}^{3}}<0,$
$\Gamma_{1}$ $=\frac{1}{6\,r_{\rm isco}^{6}}\Bigl[15q^{4}+r_{\rm isco}^{4}+352\pi P\,r_{\rm isco}^{6}+4800\pi^{2}P^{2}r_{\rm isco}^{8}$
$\quad+\,r_{\rm isco}^{2}\sqrt{g}-120\pi P\,r_{\rm isco}^{4}\sqrt{g}$
$\quad+\,q^{2}\!\left(48\pi P\,r_{\rm isco}^{4}-9\sqrt{g}\right)\Bigr]>0,$ (31)
where
[…]
It is straightforward to verify that the expression of ${\mathrm{d}M}/{\mathrm{d}r_{h}}$ here is consistent with that given in Eq. (25). Using Eq. (27) together with the above results, we find
$\frac{\mathrm{d}r_{\rm ps}}{\mathrm{d}r_{h}}=\frac{3\,r_{\rm ps}^{2}\!\left(q^{2}-r_{h}^{2}-8\pi P\,r_{h}^{4}\right)}{2\left(2q^{2}-r_{\rm ps}^{2}\right)r_{h}^{2}},$ (33)
and
$\frac{\mathrm{d}r_{\rm isco}}{\mathrm{d}r_{h}}=-\,6\,(q^{2}-r_{h}^{2}-8\pi P\,r_{h}^{4})\,r_{\rm isco}^{13}\,\sqrt{g}\;/\\ \Bigl[r_{h}^{2}\bigl(15q^{4}r_{\rm isco}^{11}+r_{\rm isco}^{15}+352\pi P\,r_{\rm isco}^{17}+4800\pi^{2}P^{2}r_{\rm isco}^{19}\\ +48\pi P\,q^{2}\,r_{\rm isco}^{15}+(r_{\rm isco}^{2}-120\pi P\,r_{\rm isco}^{4}-9q^{2})\,r_{\rm isco}^{11}\,\sqrt{g}\bigr)\Bigr].$ (34)
In the Sch-AdS limit ($q=0$), the above relations reduce to
$\frac{\mathrm{d}r_{\rm ps}}{\mathrm{d}r_{h}}=\frac{3}{2}+12\pi Pr_{h}^{2},$ (35)
[…]
where $j=1+176\pi P\,r_{\mathrm{isco}}^{2}+1600\pi^{2}P^{2}r_{\mathrm{isco}}^{4}$. Further taking the Schwarzschild limit ($q=0$, $P=0$), one obtains
$\frac{\mathrm{d}r_{\rm ps}}{\mathrm{d}r_{h}}=\frac{3}{2},\qquad\frac{\mathrm{d}r_{\rm isco}}{\mathrm{d}r_{h}}=3,$ (37)
which are consistent with the well-known relations $r_{\rm ps}=3M$, $r_{\rm isco}=6M$, and $r_{h}=2M$.
In Figs. 2 and 3, we present the numerically computed profiles of $r_{\mathrm{ps}}(r_{h})$ and $r_{\mathrm{isco}}(r_{h})$ for the four-dimensional case. As shown, $r_{\mathrm{c}}(r_{h})$ remains globally monotonic for the Sch-AdS black hole, whereas for the RN-AdS black hole it first decreases and then increases, with the extremum occurring precisely at the point where $\mathrm{d}M/\mathrm{d}r_{h}=\kappa=0$. Truncating the curve at this point, $r_{\mathrm{c}}(r_{h})$ retains its local monotonicity. These numerical results are in complete agreement with Eqs. (33)–(36).
Figure 2: Behavior of $r_{\rm ps}(r_{h})$ along isobaric paths in $d=4$. For the Sch-AdS case, the pressure is likewise set to the value $P=0.8\,P_{\mathrm{c}}$ corresponding to the $q=1$ reference point. The same choice is adopted for subsequent figures and will not be repeated.
Figure 3: Behavior of $r_{\rm isco}(r_{h})$ along isobaric paths in $d=4$
III.2 Isothermal Process
The isothermal process corresponds to the $P$-$V$ ensemble commonly employed in black hole thermodynamics. Here the only variable parameters are $M$ and $P$, which are related through the isothermal constraint $T(r_{h};M,P)=T_{0}$. Therefore, we take $\lambda=(M,P;\lambda_{\mathrm{fixed}})$. Substituting this into the criterion (21) and choosing $r_{h}$ as the path parameter gives (see Appendix C for justification)
$\frac{\mathrm{d}r_{\rm c}}{\mathrm{d}r_{h}}=-\,\frac{1}{\Gamma_{\epsilon}}\!\left(\frac{\mathrm{d}M}{\mathrm{d}r_{h}}\,\partial_{M}\Phi_{\epsilon}+\frac{\mathrm{d}P}{\mathrm{d}r_{h}}\,\partial_{P}\Phi_{\epsilon}\right).$ (38)
In general, this relation has no simple closed form. However, for photons ($\epsilon=0$), many black-hole metrics satisfy $\partial_{P}\Phi_{0}=0$, for example when the pressure enters the metric only through a term proportional to $Pr^{2}$. In such cases, the criterion (38) reduces to
$\frac{\mathrm{d}r_{\rm ps}}{\mathrm{d}r_{h}}=-\,\frac{1}{\Gamma_{0}}\,\frac{\mathrm{d}M}{\mathrm{d}r_{h}}\partial_{M}\Phi_{0}.$ (39)
This result is consistent with the criterion for the isobaric process, but differs in that under the isobaric constraint the first law of thermodynamics requires ${\mathrm{d}M}/{\mathrm{d}r_{h}}\neq 0$, ensuring that $r_{\mathrm{ps}}(r_{h})$ remains monotonic. In contrast, under the isothermal constraint the first law takes the form
$\mathrm{d}M=T_{0}\,\mathrm{d}S+V\,\mathrm{d}P,$ (40)
where $V$ is the thermodynamic volume. Dividing both sides by $\mathrm{d}r_{h}$ yields
$\frac{\mathrm{d}M}{\mathrm{d}r_{h}}=T_{0}\,S^{\prime}(r_{h})+V\,\frac{\mathrm{d}P}{\mathrm{d}r_{h}},$ (41)
indicating that ${\mathrm{d}M}/{\mathrm{d}r_{h}}$ may vanish, which in turn allows $r_{\mathrm{ps}}(r_{h})$ to become non-monotonic. This behavior is physically intuitive: when ${\mathrm{d}M}/{\mathrm{d}r_{h}}=0$, one must have
$\frac{\mathrm{d}P}{\mathrm{d}r_{h}}=-\frac{T_{0}}{V}\,S^{\prime}(r_{h})<0.$ (42)
This means that, near an extremum of $M(r_{h})$, the same $M$ corresponds to two distinct $P$. As $r_{\rm ps}$ is $P$-independent (since $\partial_{P}\Phi_{0}=0$) while $r_{h}$ depends on $P$, consequently a single $r_{\rm ps}$ maps to two different $r_{h}$, and $r_{\rm ps}(r_{h})$ is necessarily non-monotonic. Next, we apply the criterion (38) to the metric (23).
III.2.1 Photon sphere ($\epsilon=0$)
For the metric (23) one obtains
$\frac{\mathrm{d}M}{\mathrm{d}r_{h}}=\frac{1}{(d-1)\,c_{d}}\Big[$  $2(d-3)\,r_{h}^{\,d-4}+4(d-2)\,\pi T_{0}\,r_{h}^{\,d-3}$ (43)
$-2(d-2)(d-3)\,q^{2}\,r_{h}^{\,2-d}\Big],$
and
$\Gamma_{0}=\frac{2(d-3)}{r_{\rm ps}}\Big[(d-2)\,q^{2}\,r_{\rm ps}^{\,6-2d}-1\Big].$ (44)
Moreover, $\partial_{P}\Phi_{0}=0$, and $\partial_{M}\Phi_{0}$ agrees with Eq. (28).
We now analyze the sign of ${\mathrm{d}M}/{\mathrm{d}r_{h}}$. For $q=0$ one immediately has ${\mathrm{d}M}/{\mathrm{d}r_{h}}>0$. For $q\neq 0$, differentiating ${\mathrm{d}M}/{\mathrm{d}r_{h}}$ gives
$\frac{\mathrm{d}^{2}M}{\mathrm{d}r_{h}^{2}}=\frac{1}{(d-1)c_{d}}$ [ $\Big[2(d-3)(d-4)\,r_{h}^{\,d-5}$ (45)
$+4(d-2)(d-3)\pi T_{0}\,r_{h}^{\,d-4}$
$+2(d-2)^{2}(d-3)q^{2}\,r_{h}^{\,1-d}\Big]>0,$
so ${\mathrm{d}M}/{\mathrm{d}r_{h}}$ is strictly increasing in $r_{h}$. Meanwhile, From Eq. (43),
$\lim_{r_{h}\to 0^{+}}\frac{\mathrm{d}M}{\mathrm{d}r_{h}}=-\infty,\qquad\lim_{r_{h}\to\infty}\frac{\mathrm{d}M}{\mathrm{d}r_{h}}=+\infty,$ (46)
hence there exists a unique $r_{h0}$ such that $\bigl.{\mathrm{d}M}/{\mathrm{d}r_{h}}\bigr|_{r_{h0}}=0$. By the criterion (39), it follows that:
  • $q=0$: $r_{\rm ps}(r_{h})$ is monotonically increasing;
  • $q\neq 0$: $r_{\rm ps}(r_{h})$ is non-monotonic. It decreases for small $r_{h}$, reaches a single minimum at $r_{h0}$, and then increases.
This result reveals that the presence of electric charge $q$ breaks the monotonic behavior of $r_{\mathrm{ps}}(r_{h})$, making it globally non-monotonic–an aspect that has been omitted in previous studies. The explicit expression for ${dr_{\mathrm{ps}}}/{dr_{h}}$ reads
$\frac{\mathrm{d}r_{\rm ps}}{\mathrm{d}r_{h}}$ $=\Big[(d-3)(2-d)\,q^{2}\,r_{h}^{6}$ (47)
$\qquad+\big(d-3+2(d-2)\pi T_{0}\,r_{h}\big)r_{h}^{2d}\Big]\,r_{\rm ps}^{\,4+d}\;/$
$\quad\Big[(d-3)\big((2-d)\,q^{2}\,r_{\rm ps}^{6}+r_{\rm ps}^{2d}\big)\,r_{h}^{\,4+d}\Big],$
For $d=4$ (the RN-AdS case), this reduces to
$\frac{\mathrm{d}r_{\rm ps}}{\mathrm{d}r_{h}}=\frac{r_{\rm ps}^{2}\,\left[2q^{2}-r_{h}^{2}\,(1+4\pi T_{0}\,r_{h})\right]}{(2q^{2}-r_{\rm ps}^{2})\,r_{h}^{2}},$ (48)
and, in the Sch-AdS limit ($q=0$), further simplifies to
$\frac{\mathrm{d}r_{\rm ps}}{\mathrm{d}r_{h}}=1+4\pi T_{0}\,r_{h}.$ (49)
This coincides with the standard Sch-AdS result, from which ${\mathrm{d}r_{\rm ps}}/{\mathrm{d}r_{h}}>0$ is immediately evident.
Figure 4 presents the numerically computed profiles of $r_{\mathrm{ps}}(r_{h})$. For the Sch-AdS black hole, $r_{\mathrm{ps}}(r_{h})$ begins at $r_{h,\min}=(d-3)/(4\pi T_{0})$ and increases monotonically thereafter, remaining globally monotonic in full agreement with Eq. (49). In contrast, the RN-AdS black hole exhibits the expected non-monotonic behavior: although ${\mathrm{d}M}/{\mathrm{d}r_{h}}=0$ at the extremum, the surface gravity remains positive ($\kappa=2\pi T_{0}>0$), indicating that the black hole is nonextremal and the curve need not be truncated.
The breakdown of monotonicity naturally raises an important question: does it affect the discontinuous jump of $r_{\mathrm{ps}}$ across the phase transition, potentially obscuring its optical signature?
The answer is: possibly. If the small and large black hole radii, $r_{h,s}$ and $r_{h,l}$, lie on opposite sides of the extremum of $r_{\mathrm{ps}}(r_{h})$, the amplitude of the discontinuous change in $r_{\mathrm{ps}}$ may be reduced, requiring higher observational resolution to detect it. In the extreme case where $r_{\mathrm{ps}}(r_{h,s})=r_{\mathrm{ps}}(r_{h,l})$, the discontinuity would vanish entirely, rendering the phase transition optically unobservable.
From a thermodynamic perspective, however, this is not necessarily a drawback. As shown earlier, $r_{\mathrm{c}}(r_{h})$ remains monotonic in the isobaric process, while $r_{\mathrm{ps}}(r_{h})$ becomes non-monotonic in the isothermal one. Such a contrast in optical behavior between the two ensembles offers a potential diagnostic for identifying the underlying thermodynamic process of a black hole– a perspective that provides a new avenue for probing black hole thermodynamics, to be explored in future work.
Another possibility is that $r_{h,s}$ and $r_{h,l}$ lie on the same side of the extremum of $r_{\mathrm{ps}}(r_{h})$, so that $r_{\mathrm{ps}}(r_{h})$ remains locally monotonic throughout the phase-transition region. This scenario indeed occurs for the RN-AdS black hole studied here, and in Appendix F we provide an explicit proof for the $d=4$ case.
Figure 5 shows the numerically computed profiles of $P(r_{h})$ and $P(r_{\mathrm{ps}})$ at fixed parameters, where the characteristic S-shaped behavior of the phase transition is fully preserved. In this configuration, one can observe both the global non-monotonicity of $r_{\mathrm{ps}}(r_{h})$ and the discontinuous jump of $r_{\mathrm{ps}}$ across the transition. The coexistence of these two optical signatures provides an especially powerful means for extracting thermodynamic information from black holes.
Figure 4: Behavior of $r_{\rm ps}(r_{h})$ along isothermal paths in $d=4$. The marked point on the curve corresponds to $P=0$ at $r_{h_{\min}}=(d-3)/(4\pi T_{0})$, and the unphysical portion with $P<0$ is truncated. For the Sch-AdS case, the temperature is likewise set to the value $T=0.8\,T_{\mathrm{c}}$ corresponding to the $q=1$ reference point. The same choice is adopted for subsequent figures and will not be repeated.
Figure 5: Isothermal profiles of $P(r_{h})$ and $P(r_{\rm ps})$ for $d=4$ and $q=1$.
III.2.2 ISCO ($\epsilon=1$)
The metric (23) yields
$\frac{\mathrm{d}P}{\mathrm{d}r_{h}}=-\frac{(d-2)}{8\pi}\Bigl[$ $(d-2)(d-3)\,q^{2}\,r_{h}^{\,3-2d}$ (50)
$+\bigl(3-d+2\pi T_{0}r_{h}\bigr)\,r_{h}^{-3}\Bigr],$
and
$\partial_{P}\Phi_{1}=\frac{16\pi\,r_{\mathrm{isco}}}{(d-2)(d-1)}$ [ $\Bigl[8+4(d-1)(d-2)q^{2}r_{\mathrm{isco}}^{6-2d}$ (51)
$-c_{d}(d-1)(d+1)M\,r_{\mathrm{isco}}^{3-d}\Bigr],$
where $\partial_{P}\Phi_{1}>0$ (see Appendix D for proof). Since van der Waals-type behavior occurs only for charged black holes ($q\neq 0$), we analyze the RN-AdS and Sch-AdS ($q=0$) cases separately. Moreover, because the explicit expression for ${\mathrm{d}r_{\mathrm{c}}}/{\mathrm{d}r_{h}}$ is algebraically cumbersome, we focus on whether it changes sign rather than displaying the full formula.
(i) Sch-AdS ($q=0$)
For $q=0$, Eq. (50) reduces to
$\frac{\mathrm{d}P}{\mathrm{d}r_{h}}=-\,\frac{(d-2)\,\bigl(3-d+2\pi T_{0}r_{h}\bigr)}{8\pi\,r_{h}^{3}}.$ (52)
It has a single zero at $r_{h1}={(d-3)}/{2\pi T_{0}}$, with ${\mathrm{d}P}/{\mathrm{d}r_{h}}>0$ for $r_{h}<r_{h1}$ and ${\mathrm{d}P}/{\mathrm{d}r_{h}}<0$ for $r_{h}>r_{h1}$. Using the criterion (38), together with ${\mathrm{d}M}/{\mathrm{d}r_{h}}\,\partial_{M}\Phi_{1}<0$, $\Gamma_{1}>0$, and $\partial_{P}\Phi_{1}>0$, one thus expects $r_{\rm isco}(r_{h})$ to first decrease and then increase. To make this more explicit, we rewrite the criterion (38) as
$\frac{\mathrm{d}r_{\rm c}}{\mathrm{d}r_{h}}=-\,\frac{1}{\Gamma_{\epsilon}}\,\delta(r_{\rm isco},r_{h},d,q,T_{0}).$ (53)
For sufficiently large $r_{h}$ one has ${\mathrm{d}P}/{\mathrm{d}r_{h}}<0$, which forces $\delta<0$. It remains to verify that $\delta>0$ for small $r_{h}$, ensuring a single zero crossing. In this spirit we find
$\begin{cases}\delta\bigl(3r_{h_{\min}},\,r_{h_{\min}},\,4,\,0,\,T_{0}\bigr)=\dfrac{1232\,\pi^{2}T_{0}^{2}}{27}>0,&d=4,\\[6.0pt] \delta\bigl(+\infty,\,r_{h_{\min}},\,d,\,0,\,T_{0}\bigr)=\dfrac{512\,\pi^{3}T_{0}^{3}\,r_{\rm isco}}{(d-3)^{2}(d-1)}>0,&d>4,\end{cases}$ (54)
where $r_{h_{\min}}=(d-3)/(4\pi T_{0})$ is defined by $P(r_{h_{\min}})=0$. For $d=4$ and $P=0$, the metric (23) reduces to the Schwarzschild case, giving $r_{\rm isco}=3\,r_{h_{\min}}$. For $d>4$ and $P\to 0$, the higher-dimensional Schwarzschild black hole has no ISCO, hence $r_{\rm isco}\to+\infty$. Equation (54) then confirms that $\delta$ indeed crosses zero. Consequently, under the isothermal constraint, the function $r_{\rm isco}(r_{h})$ is non-monotonic.
(ii) RN-AdS ($q\neq 0$)
Since the RN-AdS black hole exhibits a van der Waals–like phase transition, one has
$\frac{\mathrm{d}P}{\mathrm{d}r_{h}}\leq 0\quad\text{for}\,T\geq T_{c},$ (55)
where the equality holds precisely at the critical point $(T_{c},P_{c},r_{hc})$. According to the criterion (38), when $r_{h}\geq r_{h0}$ (where $r_{h0}$ is the zero of $\mathrm{d}M/\mathrm{d}r_{h}$), it follows that $\delta<0$ and hence $\mathrm{d}r_{\mathrm{isco}}/\mathrm{d}r_{h}>0$, so $r_{\mathrm{isco}}(r_{h})$ increases monotonically. For $r_{h}<r_{h0}$, however, $\delta$ may become positive. To determine this behavior, we examine the limit $r_{h}\to 0^{+}$.
From Eq. (41), one obtains
$\lim_{r_{h}\to 0^{+}}\frac{\mathrm{d}M}{\mathrm{d}r_{h}}=V\,\frac{\mathrm{d}P}{\mathrm{d}r_{h}},$ (56)
which substituted into the criterion (38) gives
$\delta=\frac{\mathrm{d}P}{\mathrm{d}r_{h}}\!\left(V\,\partial_{M}\Phi_{1}+\partial_{P}\Phi_{1}\right).$ (57)
Meanwhile, from Appendix E, as $r_{h}\to 0^{+}$ one has
$V\,\partial_{M}\Phi_{1}\ \longrightarrow\ -\infty,\qquad\partial_{P}\Phi_{1}\ \longrightarrow\ 0.$ (58)
Hence,
$\delta=\frac{\mathrm{d}P}{\mathrm{d}r_{h}}\Bigl(V\,\partial_{M}\Phi_{1}+\partial_{P}\Phi_{1}\Bigr)>0,\quad\Rightarrow\quad\frac{\mathrm{d}r_{\mathrm{isco}}}{\mathrm{d}r_{h}}<0.$ (59)
This demonstrates that $r_{\mathrm{isco}}(r_{h})$ is non-monotonic over the full domain of $r_{h}$.
The same conclusion holds for $T<T_{c}$. Along both the small–black–hole and large–black–hole stable branches, one still has $\mathrm{d}P/\mathrm{d}r_{h}<0$, and the leading-order behavior as $r_{h}\to 0^{+}$ remains identical to the $T\geq T_{c}$ case above. The only difference is that for $T<T_{c}$ an intermediate spinodal (unstable) branch appears, on which $\mathrm{d}P/\mathrm{d}r_{h}>0$. Across this segment, the sign of $\delta$ may reverse, thereby altering the local monotonicity of $r_{\mathrm{isco}}(r_{h})$. In all cases, however, under the isothermal constraint $r_{\mathrm{isco}}(r_{h})$ is globally non-monotonic.
Figure 6 presents the numerically computed profiles of $r_{\mathrm{isco}}(r_{h})$. It can be seen that $r_{\mathrm{isco}}(r_{h})$ is non-monotonic for both charged and uncharged cases. The distinction lies in that, for the uncharged (Sch-AdS) case, $r_{\mathrm{isco}}(r_{h})$ starts at $r_{h_{\min}}=(d-3)/(4\pi T_{0})$ where $P(r_{h})=0$, whereas for the charged (RN-AdS) black hole, $P(r_{h})>0$ holds throughout the entire domain.
Appendix F likewise shows that $r_{h,s}$ and $r_{h,l}$ lie on the same side of the extremum of $r_{\mathrm{isco}}(r_{h})$. Consequently, $P(r_{\mathrm{isco}})$ preserves the full local SS–shaped structure of $P(r_{h})$ across the phase transition. An explicit example for fixed parameters is shown in Fig. 7.
Figure 6: Behavior of $r_{\rm isco}(r_{h})$ along isothermal paths in $d=4$.
Figure 7: Isothermal profiles of $P(r_{h})$ and $P(r_{\rm isco})$ for $d=4$ and $q=1$.
It is worth noting that, according to the first law of thermodynamics, a non-monotonic $M(r_{h})$ implies that the same mass $M$ can correspond to two distinct pressures $P$. Since both $r_{\mathrm{isco}}$ and $r_{h}$ depend on $P$, one might expect $r_{\mathrm{isco}}(r_{h})$ to be less prone to non-monotonic behavior. However, our results show that non-monotonicity indeed occurs, even for the Sch-AdS black hole.
To better understand this, we examine the criterion (38). Because $\partial_{M}\Phi_{1}<0$ and $\partial_{P}\Phi_{1}>0$, $r_{\mathrm{isco}}(r_{h})$ is guaranteed to be monotonic as long as ${\mathrm{d}M}/{\mathrm{d}r_{h}}$ and ${\mathrm{d}P}/{\mathrm{d}r_{h}}$ have opposite signs. Yet the first law,
$\frac{\mathrm{d}M}{\mathrm{d}r_{h}}=T_{0}\,S^{\prime}(r_{h})+V\,\frac{\mathrm{d}P}{\mathrm{d}r_{h}},$ (60)
requires that ${\mathrm{d}M}/{\mathrm{d}r_{h}}$ and ${\mathrm{d}P}/{\mathrm{d}r_{h}}$ share the same sign in certain ranges of $r_{h}$.
For the $d$-dimensional RN-AdS black hole, ${\mathrm{d}M}/{\mathrm{d}r_{h}}$ is non-monotonic and can take negative values. Since the first law demands ${\mathrm{d}P}/{\mathrm{d}r_{h}}<0$, the two derivatives become of the same sign, allowing the criterion (38) to cross zero. For the $d$-dimensional Sch-AdS black hole, comparing
$T_{0}S^{\prime}(r_{h})=\frac{4\pi}{c_{d}}\,(d-2)\,T_{0}\,r_{h}^{\,d-3},$ (61)
with
$\frac{\mathrm{d}M}{\mathrm{d}r_{h}}=\frac{2(d-3)\,r_{h}^{\,d-4}}{(d-1)c_{d}}+\frac{4\pi(d-2)T_{0}\,r_{h}^{\,d-3}}{(d-1)c_{d}},$ (62)
immediately reveals their relative size in the small-horizon regime. For instance, at the minimal value $r_{h_{\min}}=(d-3)/(4\pi T_{0})$, one finds
$\frac{\mathrm{d}M}{\mathrm{d}r_{h}}$ $=\frac{d}{d-1}\frac{(4\pi)^{4-d}T_{0}}{c_{d}}(\frac{d-3}{T_{0}})^{d-3},$ (63)
$T_{0}S^{\prime}(r_{h})$ $=(d-2)\frac{(4\pi)^{4-d}T_{0}}{c_{d}}(\frac{d-3}{T_{0}})^{d-3}.$ (64)
It is clear that
$\frac{\mathrm{d}M}{\mathrm{d}r_{h}}<T_{0}S^{\prime}(r_{h}),$
which, through the first law $\frac{\mathrm{d}M}{\mathrm{d}r_{h}}=T_{0}S^{\prime}(r_{h})+V\,\frac{\mathrm{d}P}{\mathrm{d}r_{h}},$ requires $\frac{\mathrm{d}P}{\mathrm{d}r_{h}}>0$. Thus ${\mathrm{d}M}/{\mathrm{d}r_{h}}$ and $\mathrm{d}P/\mathrm{d}r_{h}$ acquire the same sign, allowing the criterion (38) to cross zero.
IV Critical-exponent invariance
Critical exponents quantify the universal scaling near the black-hole critical point. They are conventionally defined with the horizon radius $r_{h}$, yet the interchangeability of $r_{\rm c}$ with $r_{h}$ suggests that one may equally define them using the circular-orbit radius $r_{\rm c}$. In this section we prove that, for both isobaric and isothermal constraints, the mapping $r_{\rm c}(r_{h})$ is locally monotonic at criticality, so the replacement $r_{h}\mapsto r_{\rm c}$ amounts to a smooth linear reparametrization in a neighborhood of the critical point. Consequently, critical exponent defined with $r_{\rm c}$ coincides with its counterpart defined with $r_{h}$.
The order-parameter exponent $\beta$ is defined from the coexistence gap of the horizon radius for $T<T_{c}$:
$\Delta r_{h}:=r_{h,l}-r_{h,s}\sim|t|^{\beta},\qquad t:=\frac{T-T_{c}}{T_{c}}\to 0^{-},$ (65)
with the mean-field value $\beta=1/2$. We will show that the circular-orbit radius inherits the same scaling, namely
$\Delta r_{c}\;:=\;r_{c}\!\big(r_{h,l}\big)-r_{c}\!\big(r_{h,s}\big)\;\sim\;|t|^{1/2}.$ (66)
We first perform a Taylor expansion of $P(r_{h},T)$ about the critical point $(r_{h,c},T_{c})$ up to the first nonvanishing orders allowed by the criticality conditions $\bigl.(\partial_{r_{h}}P)\bigr|_{c}=0$ and $\bigl.(\partial_{r_{h}}^{2}P)\bigr|_{c}=0$:
$P(r_{h},T)-P_{c}$ $=\left.(\partial_{T}P)\right|_{c}\,\Delta T+\left.(\partial_{T}\partial_{r_{h}}P)\right|_{c}\,\Delta T\,\Delta r$ (67)
$+\frac{1}{2}\left.(\partial_{r_{h}}^{2}\partial_{T}P)\right|_{c}\,\Delta T\,(\Delta r)^{2}$
$+\frac{1}{6}[\left.(\partial_{r_{h}}^{3}P)\right|_{c}+\left.(\partial_{T}\partial_{r_{h}}^{3}P)\right|_{c}\,\Delta T]\,(\Delta r)^{3}$
$+\mathcal{O}\!\big(\Delta T^{2},\,(\Delta r)^{4}\big),$
where $\Delta r:=r_{h}-r_{h,c}$ and $\Delta T:=T-T_{c}$. At fixed temperature $T$, differentiating with respect to $r_{h}$ gives
$\partial_{r_{h}}P(r_{h},T)$ $=\left.(\partial_{T}\partial_{r_{h}}P)\right|_{c}\,\Delta T$ (68)
$+\left.(\partial_{r_{h}}^{2}\partial_{T}P)\right|_{c}\,\Delta T\,(\Delta r)$
$+\frac{1}{2}[\left.(\partial_{r_{h}}^{3}P)\right|_{c}+\left.(\partial_{T}\partial_{r_{h}}^{3}P)\right|_{c}\,\Delta T]\,(\Delta r)^{2}$
$+\mathcal{O}\!\big(\Delta T^{2},\,(\Delta r)^{4}\big),$
On the coexistence line for $T<T_{c}$, the Eq. (65) gives
$\Delta r=(r_{h,l}-r_{h,c})-(r_{h,s}-r_{h,c})=\Delta r_{h}\sim|t|^{1/2},$ (69)
so each power counts as $\Delta r^{n}\sim|t|^{n/2}$. Substituting this into (67) and keeping the leading orders in $|t|$ yields
$\partial_{r_{h}}P(r_{h},T)$ (70)
$=\left.(\partial_{T}\partial_{r_{h}}P)\right|_{c}\,T_{c}\,t+\frac{1}{2}\left.(\partial_{r_{h}}^{3}P)\right|_{c}\,|t|+\mathcal{O}\!\big(|t|^{3/2}\big),$
Thus $\bigl(\partial P/\partial r_{h}\bigr)_{T}=\mathcal{O}(|t|)\to 0$ as $t\to 0^{-}$. Along an isotherm, our criterion
$\frac{\mathrm{d}r_{\rm c}}{\mathrm{d}r_{h}}=-\,\frac{1}{\Gamma_{\epsilon}}\!\left(\frac{\mathrm{d}M}{\mathrm{d}r_{h}}\,\partial_{M}\Phi_{\epsilon}+\frac{\mathrm{d}P}{\mathrm{d}r_{h}}\,\partial_{P}\Phi_{\epsilon}\right)$ (71)
combines with the first law
$\frac{\mathrm{d}M}{\mathrm{d}r_{h}}=T\,S^{\prime}(r_{h})+V\,\frac{\mathrm{d}P}{\mathrm{d}r_{h}}$ (72)
to give, as $t\to 0^{-}$,
$\frac{\mathrm{d}M}{\mathrm{d}r_{h}}=T_{c}\,S^{\prime}(r_{h,c})+\mathcal{O}(|t|),\qquad\frac{\mathrm{d}P}{\mathrm{d}r_{h}}=\mathcal{O}(|t|).$ (73)
Since $S^{\prime}(r_{h,c})>0$, $V>0$, and $\Gamma_{\epsilon}\neq 0$, $\partial_{M}\Phi_{\epsilon}$ and $\partial_{P}\Phi_{\epsilon}$ are finite at the critical point, we obtain the fixed-sign leading term
$\frac{\mathrm{d}r_{\rm c}}{\mathrm{d}r_{h}}=-\,\frac{T_{c}\,S^{\prime}(r_{h,c})}{\Gamma_{\epsilon}}\,\partial_{M}\Phi_{\epsilon}\ +\ \mathcal{O}(|t|),$ (74)
so $r_{c}(r_{h})$ is strictly monotonic for sufficiently small $|t|$. (In the photon case, $\partial_{P}\Phi_{0}=0$ further collapses the subleading $\mathcal{O}(|t|)$ contribution.)
Monotonicity (74) implies the first-order linearization at $r_{h,c}$,
$r_{\rm c}(r_{h})=r_{{\rm c},c}+C\,\bigl(r_{h}-r_{h,c}\bigr)+o\!\bigl(|r_{h}-r_{h,c}|\bigr),$ (75)
where $C:=\left.{dr_{c}}/{dr_{h}}\right|_{c}\neq 0$. Therefore on the coexistence line
$\Delta r_{c}=r_{c,l}-r_{c,s}=C\,\Delta r_{h}+o(\Delta r_{h})\sim|t|^{\beta},$ (76)
and the critical exponent is unchanged:
$\beta_{(r_{c})}=\beta_{(r_{h})}=\tfrac{1}{2}\quad\text{(mean-field)}.$ (77)
For the isobaric process, $r_{c}(r_{h})$ is globally monotonic, so the above conclusion applies verbatim. In summary, under either isobaric or isothermal constraints, the reparameterization $r_{h}\!\mapsto\!r_{c}$ does not change the critical exponents. This is consistent with the numerical findings of Wei et al. in the isobaric photon-sphere case [12], for which we have provided an analytic proof.
V Discussion
In this work, we derived the general criterion (21), applicable to any parameter path in spherically symmetric black holes, which clarifies how circular orbits encode thermodynamic behavior. The criterion reveals that the monotonic relation between the circular-orbit radius $r_{\mathrm{c}}$ and the horizon radius $r_{h}$ is influenced by two factors: the intrinsic properties of the black hole and its circular orbits, and the parameter path chosen in the extended thermodynamic space. While the former never breaks monotonicity, the latter plays a decisive role. Motivated by this, we investigated two thermodynamically significant paths–the isobaric and isothermal processes–using the $d$-dimensional RN-AdS black hole as a test case.
Our analysis demonstrates that the first law of thermodynamics directly constrains the monotonicity of $r_{\mathrm{c}}(r_{h})$. Under the isobaric constraint, the first law enforces $\mathrm{d}M/\mathrm{d}r_{h}\neq 0$, ensuring that $M(r_{h})$ and hence $r_{\mathrm{c}}(r_{h})$ remain monotonic. In contrast, under the isothermal constraint the first law allows $\mathrm{d}M/\mathrm{d}r_{h}=0$, which decouples the extrema of $M(r_{h})$ and $P(r_{h})$ so that a single mass $M$ corresponds to two distinct pressures $P$. This inevitably makes $r_{\mathrm{ps}}(r_{h})$ non-monotonic, since the photon-sphere radius $r_{\mathrm{ps}}$ is $P$-independent whereas $r_{h}$ depends on $P$. An exception occurs in the Sch-AdS limit ($q=0$), where $M(r_{h})$ remains monotonic; the emergence of electric charge $q$ thus breaks this monotonicity and, in turn, that of $r_{\mathrm{ps}}(r_{h})$.
Interestingly, for the ISCO case, both $r_{\mathrm{isco}}$ and $r_{h}$ depend explicitly on $P$, which might seem to favor monotonicity. However, for the RN-AdS family we find that $r_{\mathrm{isco}}(r_{h})$ is non-monotonic for all $d$, including the Sch-AdS limit. This arises because the first law requires $\mathrm{d}M/\mathrm{d}r_{h}$ and $\mathrm{d}P/\mathrm{d}r_{h}$ to share the same sign in certain domains, making the prefactor in the criterion (38) change sign and thus breaking monotonicity. Overall, our results indicate that $r_{\mathrm{c}}(r_{h})$ tends to be monotonic under isobaric evolution but non-monotonic under isothermal evolution.
It is worth noting that this non-monotonicity does not spoil the monotonic behavior of $r_{\rm c}(r_{h})$ in the local phase-transition region, nor does it affect its behavior near the critical point. As a result, the jump in $r_{\rm c}$ remains a reliable signature of the transition, and the critical exponents defined through $r_{\rm c}$ coincide with those obtained from $r_{h}$. These features are closely tied to the black hole first law and are expected to extend to more general classes of black holes.
These results fill an important gap in the study of the non-monotonic behavior of $r_{\mathrm{c}}(r_{h})$ and provide a more robust foundation for identifying black hole phase transitions through optical observables. They also offer an additional avenue for probing black hole thermodynamics: the monotonicity (or lack thereof) of $r_{\mathrm{c}}(r_{h})$ can be used to distinguish whether a black hole evolves along an isobaric or an isothermal process.
This idea is physically well motivated. By the entropy-increase law, the horizon radius $r_{h}$ must grow with time, and if $r_{\mathrm{c}}(r_{h})$ is non-monotonic, the resulting optical appearance of the black hole will exhibit correspondingly distinct behavior. Moreover, the influence of black hole thermodynamics on the monotonicity of $r_{\mathrm{c}}(r_{h})$ is itself a noteworthy feature deserving further investigation.
Acknowledgements.
This work is supported in part by NSFC Grants No. 12165005 and No. 11961131013.
Appendix A Preservation of Functional Behavior under Monotonic Reparameterization
Suppose that $F(x)$ is a differentiable function defined on an interval $I\subset\mathbb{R}$, and let $x_{m}$ denote its $m$-th zero, so that $F(x_{m})=0$. Now introduce a new variable $y$ and assume a smooth, strictly monotonic mapping
$x=x(y),\qquad y\in J\subset\mathbb{R},$ (78)
with ${\mathrm{d}x}/{\mathrm{d}y}\neq 0$ for all $y\in J$. We then reparameterize $F(x)$ as
$\widetilde{F}(y)=F\bigl(x(y)\bigr).$ (79)
Applying the chain rule gives
$\frac{\mathrm{d}\widetilde{F}}{\mathrm{d}y}=\frac{\mathrm{d}F}{\mathrm{d}x}\,\frac{\mathrm{d}x}{\mathrm{d}y}.$ (80)
Since $\mathrm{d}x/\mathrm{d}y\neq 0$, one immediately has
$\frac{\mathrm{d}\widetilde{F}}{\mathrm{d}y}=0\quad\Longleftrightarrow\quad\frac{\mathrm{d}F}{\mathrm{d}x}=0.$ (81)
Thus, the stationary points of $F$ and $\widetilde{F}$ correspond one-to-one: each stationary point $x_{m}$ of $F$ maps to a stationary point $y_{m}=x^{-1}(x_{m})$ of $\widetilde{F}$, and vice versa.
To demonstrate that the nature of the stationary points is preserved, we differentiate once more:
$\frac{\mathrm{d}^{2}\widetilde{F}}{\mathrm{d}y^{2}}=F^{\prime\prime}\bigl(x(y)\bigr)\,[x^{\prime}(y)]^{2}+F^{\prime}\bigl(x(y)\bigr)\,x^{\prime\prime}(y).$ (82)
At a stationary point $y_{m}$ where $F^{\prime}\bigl(x(y_{m})\bigr)=0$, this reduces to
$\left.\frac{\mathrm{d}^{2}\widetilde{F}}{\mathrm{d}y^{2}}\right|_{y_{m}}=F^{\prime\prime}(x_{m})\,[x^{\prime}(y_{m})]^{2}.$ (83)
Since $x^{\prime}(y_{m})]^{2}>0$, it follows that
$F^{\prime\prime}(x_{m})>0$ $\ \Longrightarrow\ \widetilde{F}^{\prime\prime}(y_{m})>0,$ (84)
$F^{\prime\prime}(x_{m})<0$ $\ \Longrightarrow\ \widetilde{F}^{\prime\prime}(y_{m})<0.$
Hence, under any smooth and monotonic reparameterization $x\mapsto y$, a local minimum (maximum) of $F$ remains a local minimum (maximum) of $\widetilde{F}$.
Away from the stationary points, Eq. (80) implies that when $\mathrm{d}x/\mathrm{d}y>0$, the signs of $\mathrm{d}\widetilde{F}/\mathrm{d}y$ and $\mathrm{d}F/\mathrm{d}x$ coincide, whereas for $\mathrm{d}x/\mathrm{d}y<0$ they are opposite. Geometrically, the former corresponds to a horizontal stretching of the original curve, while the latter involves both stretching and a reflection about the $x$-axis. In either case, the overall qualitative profile of the function remains unchanged.
If $F$ is identified with the thermodynamic temperature $T$ or pressure $P$ of a black hole, and the variables are smoothly and monotonically related as $x=r_{h}$ and $y=r_{c}$, then under the mapping $r_{h}\mapsto r_{c}$ the functions $T(r_{c})$ and $P(r_{c})$ retain their characteristic van der Waals–like $S$-shaped structure. As a result, during the small-to-large black-hole phase transition, the critical radius $r_{c}$ undergoes a finite discontinuous jump.
Appendix B Constructive counterexample in Sch-AdS demonstrating that $r_{\rm ps}(r_{h})$ can decrease
For the four-dimensional Sch-AdS black hole, the metric function reads
$f(r)=1-\frac{2M}{r}+\frac{8\pi P}{3}\,r^{2},$ (85)
where $M$ is the ADM mass and $P=-\Lambda/(8\pi)$ denotes the thermodynamic pressure. The photon-sphere condition $rf^{\prime}(r)-2f(r)=0$ gives
$r_{\rm ps}=3M,$ (86)
indicating that $r_{\rm ps}$ is independent of $P$. The event horizon radius $r_{h}$ satisfies $f(r_{h};M,P)=0$, from which we obtain
$M=\frac{r_{h}}{2}+\frac{4\pi P}{3}\,r_{h}^{3}.$ (87)
Eliminating $M$ between Eqs. (86) and (87) yields
$r_{\rm ps}(r_{h})=\frac{3r_{h}}{2}+4\pi P\,r_{h}^{3}.$ (88)
Figure 8 illustrates $r_{\rm ps}(r_{h})$ for nine representative values of $P$, without imposing any specific relation between $P$ and $M$.
Figure 8: Plots of $r_{\rm ps}(r_{h})$ for the Sch-AdS black hole. Colored solid curves denote nine different pressures $P$, and the black dashed curve shows the path-constrained case with $k<0$ and $b>0$.
Next, we impose the path constraint
$P(M)=-\frac{3k^{2}\,\bigl(b+(-1+2k)M\bigr)}{8\pi\,(b-M)^{3}},$ (89)
where $k$ is a dimensionless constant and $b$ carries dimensions of mass. Substituting Eqs. (89), (87), and (86) gives a simple linear relation,
$r_{\rm ps}(r_{h})=3\,(k\,r_{h}+b).$ (90)
In Fig. 8, this corresponds to the black dashed line, which decreases monotonically between points $A$ and $B$. Its intersections with the colored curves are uniquely determined by the constraint (89). Along this dashed segment, all physical requirements are satisfied: $M>0$, $r_{h}>0$, $P>0$, and the surface gravity remains strictly positive,
$\kappa=\frac{1}{2}\left.\partial_{r}f(r)\right|_{r_{h}}=\frac{M}{r_{h}^{2}}+\frac{8\pi Pr_{h}}{3}>0,$ (91)
ensuring non-extremality.
This simple example illustrates that the monotonicity of $r_{\rm ps}(r_{h})$ [and likewise of $r_{\rm isco}(r_{h})$] is governed not only by the functional dependence of $(M,P)$ in $f(r)$, but also by the parameter path that links $M$ and $P$. Appropriate constraints can thus lead to a non-increasing behavior of $r_{\rm ps}(r_{h})$. Applying the criterion (21) yields
$\frac{\mathrm{d}r_{\rm ps}}{\mathrm{d}r_{h}}=\frac{2\kappa}{\Gamma_{0}}\,\frac{\partial_{M}\Phi_{0}}{\partial_{M}f+\partial_{P}f\,\frac{\mathrm{d}P}{\mathrm{d}M}},$ (92)
where
$\frac{\mathrm{d}P}{\mathrm{d}M}=-\frac{3k^{2}\,\bigl(b+bk+2kM-M\bigr)}{4\pi\,(b-M)^{4}}.$ (93)
After straightforward algebra, we obtain
$\frac{\mathrm{d}r_{\rm ps}}{\mathrm{d}r_{h}}=3k,$ (94)
which agrees with Eq. (90).
Appendix C Local Existence of $M(r_{h})$ and $P(r_{h})$
Consider the two constraint equations at the horizon,
$f(r_{h};M,P)=0,\qquad T(r_{h};M,P)=T_{0},$ (95)
where both $f$ and $T$ are assumed to be smooth in $(M,P)$. Differentiating these relations and treating $r_{h}$ as the path parameter gives
$\partial_{r_{h}}f\,dr_{h}+\partial_{M}f\,dM+\partial_{P}f\,dP$ $=0,$ (96)
$\partial_{r_{h}}T\,dr_{h}+\partial_{M}T\,dM+\partial_{P}T\,dP$ $=0.$
Eliminating $\mathrm{d}P$ and $\mathrm{d}M$ in turn yields
$\frac{\mathrm{d}M}{\mathrm{d}r_{h}}$ $=-\,\frac{\partial_{r_{h}}f\,\partial_{P}T-\partial_{P}f\,\partial_{r_{h}}T}{\partial_{M}f\,\partial_{P}T-\partial_{P}f\,\partial_{M}T},$ (97)
$\frac{\mathrm{d}P}{\mathrm{d}r_{h}}$ $=-\,\frac{\partial_{r_{h}}f\,\partial_{M}T-\partial_{M}f\,\partial_{r_{h}}T}{\partial_{M}f\,\partial_{P}T-\partial_{P}f\,\partial_{M}T}.$
Defining
$\Delta:=\partial_{M}f\,\partial_{P}T-\partial_{P}f\,\partial_{M}T,$ (98)
we see that the nondegeneracy condition $\Delta\neq 0$ guarantees–by the implicit function theorem–the local existence and uniqueness of smooth functions $M(r_{h})$ and $P(r_{h})$ satisfying the two constraints in a neighborhood of a reference point. When $\Delta=0$, this parametrization in terms of $r_{h}$ breaks down and the curve must be reparametrized. For the metric (23), one finds
$\Delta=-\,\frac{4c_{d}}{d-2}\,r_{h}^{\,4-d}\neq 0.$ (99)
Appendix D Proof of the Positivity of $\partial_{P}\Phi_{1}$
$\partial_{P}\Phi_{1}=\frac{16\pi\,r_{\mathrm{isco}}}{(d-2)(d-1)}$ [ $\Bigl[8+4(d-1)(d-2)q^{2}r_{\mathrm{isco}}^{6-2d}$ (100)
$-c_{d}(d-1)(d+1)M\,r_{\mathrm{isco}}^{3-d}\Bigr].$
Since the prefactor is positive, the sign of $\partial_{P}\Phi_{1}$ is determined by the bracketed term. Let $x:=r_{\mathrm{isco}}^{\,d-3}>0$, so that
$B=8+\frac{4(d-1)(d-2)q^{2}}{x^{2}}-\frac{(d-1)(d+1)c_{d}M}{x}.$ (101)
Multiplying both sides by $x^{2}>0$ yields an upward-opening quadratic,
$\mathcal{Q}(x)=8x^{2}-(d-1)(d+1)c_{d}Mx+4(d-1)(d-2)q^{2},$ (102)
with discriminant
$\Delta_{x}=\big[(d-1)(d+1)c_{d}M\big]^{2}-128(d-1)(d-2)q^{2},$ (103)
and roots
$x_{\pm}=\frac{(d-1)(d+1)c_{d}M\pm\sqrt{\Delta_{x}}}{16},\qquad(\Delta_{x}\geq 0).$ (104)
Hence,
$\mathrm{sgn}\,\partial_{P}\Phi_{1}=\begin{cases}+,&\Delta_{x}<0\ \text{or}\ x\in(0,x_{-})\cup(x_{+},\infty),\\[2.0pt] 0,&x=x_{\pm},\\[2.0pt] -,&x\in(x_{-},x_{+}).\end{cases}$ (105)
To determine the sign of $\partial_{P}\Phi_{1}$, we first examine the sign of $\Delta_{x}$. Under the isothermal constraint one has
$T_{0}=\frac{(d-1)c_{d}M\,r_{h}^{d+3}-2(d-2)q^{2}r_{h}^{6}-2r_{h}^{2d}}{4\pi r_{h}^{1+2d}}.$ (106)
The requirement $T_{0}>0$ is equivalent to
$(d-1)c_{d}M\,r_{h}^{d+3}-2(d-2)q^{2}r_{h}^{6}-2r_{h}^{2d}>0.$ (107)
Introducing $r_{h}^{2d}=r_{h}^{6}s^{2}$ and $r_{h}^{d+3}=r_{h}^{6}s$, we define
$\mathcal{B}(s):=-2s^{2}+(d-1)c_{d}\,M\,s-2(d-2)q^{2}>0,$ (108)
which is a downward-opening quadratic function. For $\mathcal{B}(s)>0$ to hold, its discriminant must satisfy
$\Delta_{s}=\big(c_{d}(d-1)M\big)^{2}-16(d-2)q^{2}>0,$ (109)
implying
$M>\frac{4\sqrt{d-2}}{c_{d}(d-1)}\,q.$ (110)
Substituting this inequality into Eq. (103) gives
$\Delta_{x}>0.$ (111)
Hence, $\partial_{P}\Phi_{1}$ may take positive, negative, or zero values, depending on the relative magnitude of $x$ and $x_{\pm}$.
To clarify the relation between $x$ and $x_{\pm}$, we rewrite
$\Phi_{1}(r,M,q,P,d)=P\,\partial_{P}\Phi_{1}+\widetilde{\Phi}_{1}(r,M,q,d).$ (112)
Fixing $M$, $q$, and $d$ at finite values and taking the limit $P\to\infty$, the condition $\Phi_{1}(r,M,q,P,d)=0$ can be satisfied only if
$\partial_{P}\Phi_{1}\to 0.$ (113)
This shows that the quantities $x_{\pm}^{1/(d-3)}$ correspond exactly to the limiting values of $r_{\rm isco}$ as $P\to\infty$. Next, we examine the limit $P\to 0$. For $d=4$, one obtains
$\Phi_{1}=\frac{2Mr_{\mathrm{isco}}^{2}(r_{\mathrm{isco}}-6M)+18Mq^{2}r_{\mathrm{isco}}-8q^{4}}{r_{\mathrm{isco}}^{5}}.$ (114)
The analysis can be divided into two cases:
  • $q=0$: For the Schwarzschild black hole, this yields $r_{\mathrm{isco}}=6M$. In this case, the two roots are $x_{\pm}={15M}/{4}$ and 0, and clearly $x\!\mid_{P=0}=6M>x_{+}$.
  • $q\neq 0$: For the RN black hole, it is well known that $r_{\mathrm{isco}}$ attains its minimum at $M=q$, where $r_{\mathrm{isco,min}}=4M$. From Eq. (104), $x_{+}$ decreases monotonically with $q$, reaching its maximum $x_{+,{\rm max}}={15M}/{4}$ at $q=0$, so that $x_{min}\!\mid_{P=0}=4M>x_{+,{\rm max}}={15M}/{4}$.
Therefore, for both the Schwarzschild and RN black holes one always has $x\!\mid_{P=0}>x_{+}$, and consequently $\partial_{P}\Phi_{1}>0$ in this limit.
Next, consider increasing $P$ from zero in Eq. (112). Initially, one has $\Phi_{1}=\widetilde{\Phi}_{1}(r_{\mathrm{isco}}\!\mid_{P=0},M,q,d)=0$ with $\Gamma_{1}=\partial_{r}\widetilde{\Phi}_{1}>0$. Once $P$ increases, the term $\partial_{P}\Phi_{1}>0$ appears, so $\widetilde{\Phi}_{1}$ must decrease to maintain $\Phi_{1}=0$. Since $\Gamma_{1}>0$, this requires $r_{\mathrm{isco}}$ to decrease, meaning that $x$ moves toward $x_{+}$. As $P$ continues to increase, we assume continuous dependence of $r_{\mathrm{isco}}$ and $\partial_{P}\Phi_{1}$ on $P$, so that $x$ must vary continuously from $x\!\mid_{P=0}$ to $x_{+}\!\mid_{P\to\infty}$.
Consequently, for any finite $P$, the corresponding $x$ must lie within the interval $(x_{+},\,x\!\mid_{P=0})$, which guarantees $\partial_{P}\Phi_{1}>0$ throughout. From a physical viewpoint, this reflects the suppressive effect of pressure on the ISCO radius: as $P$ increases, $r_{\mathrm{isco}}$ decreases monotonically toward $x_{+}$.
For higher dimensions ($d>4$), the argument becomes even simpler. At $P=0$, Schwarzschild and RN black holes possess no ISCO ($r_{\mathrm{isco}}\to\infty$). As $P$ increases, $r_{\mathrm{isco}}$ moves inward from infinity toward $x_{+}$, again ensuring $\partial_{P}\Phi_{1}>0$. Hence, in all cases we conclude that
$\partial_{P}\Phi_{1}>0.$ (115)
Appendix E Asymptotic Relations in the Small-Horizon Limit
We now examine the asymptotic behavior of $V\,\partial_{M}\Phi_{1}$ and $\partial_{P}\Phi_{1}$ in the small-horizon limit ($r_{h}\to 0^{+}$), focusing on their leading divergent terms.
For $\partial_{P}\Phi_{1}$, since $r_{h}\to 0^{+}$ implies $P\to\infty$, one has $\partial_{P}\Phi_{1}\to 0$ (see Appendix D for details). The thermodynamic volume takes the simple form
$V=\frac{\Omega_{d-2}\,r_{h}^{\,d-1}}{d-1},$ (116)
and therefore vanishes as $V\sim r_{h}^{\,d-1}\!\to\!0$. $\partial_{M}\Phi_{1}$ reads
$\partial_{M}\Phi_{1}=3c_{d}(d-1)(d-3)q^{2}r_{\mathrm{isco}}^{8-3d}\\ -2c_{d}^{2}(d-1)(d-3)M\,r_{\mathrm{isco}}^{5-2d}\\ -c_{d}(d-3)(d-5)r_{\mathrm{isco}}^{2-d}\\ -\frac{16c_{d}(d+1)\pi P\,r_{\mathrm{isco}}^{4-d}}{d-2},$ (117)
and its asymptotic behavior can be inferred once the leading scalings of $M$, $P$, and $r_{\mathrm{isco}}$ are determined.
Under the isothermal constraint, one finds
$M(r_{h})=\frac{\bigl(1+2\pi T_{0}r_{h}\bigr)\,r_{h}^{2(d-3)}+2(d-2)\,q^{2}}{c_{d}\,(d-1)\,r_{h}^{\,d-3}},$ (118)
and
$P(r_{h})=\frac{(d-2)\!\left[\bigl(3-d+4\pi T_{0}r_{h}\bigr)\,r_{h}^{2(d-3)}+(d-3)\,q^{2}\right]}{16\pi\,r_{h}^{\,2(d-2)}}.$ (119)
In the small-horizon limit ($r_{h}\!\to\!0^{+}$), these behave as
$M\sim\frac{1}{r_{h}^{\,d-3}},\qquad P\sim\frac{1}{r_{h}^{\,2(d-2)}}.$ (120)
For the ISCO, the divergence $P\!\to\!\infty$ implies (see Appendix D)
$r_{\mathrm{isco}}^{\,d-3}\to x_{+}=\frac{(d-1)(d+1)c_{d}M+\sqrt{\Delta_{x}}}{16},$ (121)
where
$\Delta_{x}=\big[(d-1)(d+1)c_{d}M\big]^{2}-128(d-1)(d-2)q^{2}.$ (122)
Near $r_{h}\!\to\!0^{+}$, the dominant contribution arises from the mass term $M$, so that
$r_{\mathrm{isco}}\sim M^{\tfrac{1}{d-3}}\sim\frac{1}{r_{h}}.$ (123)
Substituting Eqs. (120) and (123) into Eq. (117) gives $\partial_{M}\Phi_{1}\sim-r_{h}^{-d}$, and hence
$V\,\partial_{M}\Phi_{1}\sim-\frac{1}{r_{h}}\;\longrightarrow\;-\infty.$ (124)
Appendix F Monotonicity of $r_{\rm c}(r_{h})$ in the Phase-Transition Region
F.1 Photon sphere ($\epsilon=0$)
Our proof consists of two steps. First, we demonstrate that $r_{h0}$ lies to the left of $r_{h,\min}$. Second, we show that $P(r_{h0})>P(r_{h,\max})$, where $r_{h,\min}$ and $r_{h,\max}$ denote the locations of the minimum and maximum of $P(r_{h})$, respectively. This procedure circumvents the need to determine the small-black-hole radius $r_{h,s}$ from Maxwell’s equal-area law while still guaranteeing that $r_{h0}$ lies to the left of $r_{h,s}$.
Step 1
From Eq. (41) it follows that ${\mathrm{d}P}/{\mathrm{d}r_{h}}=0$ implies ${\mathrm{d}M}/{\mathrm{d}r_{h}}>0$. Since $M(r_{h})$ first decreases and then increases, $r_{h0}$ must therefore lie to the left of both extrema, namely on the small-black-hole branch.
Step 2
To establish the result for arbitrary $q$, we eliminate $q$ by expressing $P(r_{h})$ and ${\mathrm{d}M}/{\mathrm{d}r_{h}}$ in reduced variables:
$P(\tilde{r}_{h},\tilde{T}_{0})$ $=\frac{1-6\tilde{r}_{h}^{2}+8\tilde{r}_{h}^{3}\tilde{T}_{0}}{3\tilde{r}_{h}^{4}},$ (125)
$\frac{\mathrm{d}M}{\mathrm{d}r_{h}}(\tilde{r}_{h},\tilde{T}_{0})$ $=\frac{-1+3\tilde{r}_{h}^{2}+4\tilde{r}_{h}^{3}\tilde{T}_{0}}{6\tilde{r}_{h}^{2}},$ (126)
where $\tilde{r}_{h}=r_{h}/r_{h,c}$ and $\tilde{T}_{0}=T_{0}/T_{c}$, with $r_{h,c}=\sqrt{6}\,q$ and $T_{c}=1/(3\sqrt{6}\,\pi q)$ denoting the critical values. Using these relations, we numerically determine $r_{h0}$ and $r_{h,\max}$, and plot $P(r_{h0})$ (red curve) and $P(r_{h,\max})$ (blue curve) as functions of $\tilde{T}$ in Fig. 9. The black dashed vertical line at $\tilde{T}_{0}=\sqrt{2}/2$ marks the critical value above which $P>0$; the region $\tilde{T}_{0}<\sqrt{2}/2$ is therefore unphysical and excluded from consideration. From the remaining portion of the plot it is evident that $P(r_{h0})>P(r_{h,\max})$. Combined with the result of Step 1, this demonstrates that $r_{h0}$ lies to the left of $r_{h,s}$.
To visualize this behavior, Fig. 10 shows $P(r_{h})$ (red) and ${\mathrm{d}M}/{\mathrm{d}r_{h}}$ (blue) for several $\tilde{T}_{0}$. At $\tilde{T}_{0}=0.4$, $P(r_{h0})<P(r_{h,\max})$, but $P(r_{h})<0$ and is thus unphysical. When $\tilde{T}_{0}\geq\sqrt{2}/2$, $P(r_{h})\geq 0$ and $P(r_{h0})>P(r_{h,\max})$, with $r_{h,s}$ and $r_{h,l}$ clearly located to the right of $r_{h0}$.
F.2 ISCO ($\epsilon=1$)
With the photon-sphere case established, the ISCO analysis becomes straightforward. From the criterion
$\frac{\mathrm{d}r_{\rm c}}{\mathrm{d}r_{h}}=-\,\frac{1}{\Gamma_{\epsilon}}\!\left(\frac{\mathrm{d}M}{\mathrm{d}r_{h}}\,\partial_{M}\Phi_{\epsilon}+\frac{\mathrm{d}P}{\mathrm{d}r_{h}}\,\partial_{P}\Phi_{\epsilon}\right),$ (127)
we note that, if the term $\mathrm{d}P/\mathrm{d}r_{h}\,\partial_{P}\Phi_{1}$ is momentarily ignored, the zeros of ${\mathrm{d}M}/{\mathrm{d}r_{h}}\,\partial_{M}\Phi_{1}$ coincide with those of ${\mathrm{d}M}/{\mathrm{d}r_{h}}\,\partial_{M}\Phi_{0}$, since both are controlled entirely by $\mathrm{d}M/\mathrm{d}r_{h}$. Because $\partial_{M}\Phi_{1}<0$ and $\mathrm{d}M/\mathrm{d}r_{h}$ increases monotonically, ${\mathrm{d}M}/{\mathrm{d}r_{h}}\,\partial_{M}\Phi_{1}$ is strictly decreasing.
Restoring the term $\mathrm{d}P/\mathrm{d}r_{h}\,\partial_{P}\Phi_{1}<0$ can only shift the zero of $\frac{\mathrm{d}M}{\mathrm{d}r_{h}}\,\partial_{M}\Phi_{1}+\frac{\mathrm{d}P}{\mathrm{d}r_{h}}\,\partial_{P}\Phi_{1}$further to the left. Since the zero of $\mathrm{d}M/\mathrm{d}r_{h}$ already lies to the left of $r_{h,s}$ (as shown earlier), its additional leftward shift necessarily keeps it on the small-black-hole side.
Therefore, under the isothermal constraint, although $r_{\mathrm{c}}(r_{h})$ of the four-dimensional RN-AdS black hole is globally non-monotonic, it remains locally monotonic within the phase-transition region, thereby preserving the discontinuous jump across the transition.
Figure 9: Plots of $p(r_{h0})$ (red) and $p(r_{h,\max})$ (blue) as functions of $\tilde{T}_{0}$. At $\tilde{T}_{0}=\sqrt{2}/2$, one has $p(r_{h})>0$ for all $r_{h}$.
Figure 10: Behaviors of $P(r_{h})$ and $\mathrm{d}M/\mathrm{d}r_{h}$ for three representative values of $\tilde{T}_{0}$.